OpenStudy (anonymous):

Find cot θ if csc θ = square root of seventeen divided by four and tan θ > 0. a. square root of seventeen b. one divided by four c. square root of seventeen divided by seventeen d. 17

OpenStudy (anonymous):

\[\csc(\theta)=\frac{\sqrt{17}}{4}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

find \(a\) via pythagoras then you will have it

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

@hartnn please help

hartnn (hartnn):

you need help with this question? what have you tried?

OpenStudy (anonymous):

yes